# The DSA Transparency Database: Auditing Self-reported Moderation Actions by Social Media

### 16 Dec 2023 v1

@@ Abstract
Since September 2023, the Digital Services Act (DSA) obliges large online platforms to submit detailed data on each moderation action they take within the European Union (EU) to the DSA Transparency Database. From its inception, this centralized database has sparked scholarly interest as an unprecedented and potentially unique trove of data on real-world online moderation. Here, we thoroughly analyze all 195.61M records submitted by the eight largest social media platforms in the EU during the first 60 days of the database. Specifically, we conduct a platform-wise comparative study of their: volume of moderation actions, grounds for decision, types of applied restrictions, types of moderated content, timeliness in undertaking and submitting moderation actions, and use of automation. Furthermore, we systematically cross-check the contents of the database with the platforms’ own transparency reports. Our analyses reveal that (i) the platforms adhered only in part to the philosophy and structure of the database, (ii) the structure of the database is partially inadequate for the platforms’ reporting needs, (iii) the platforms exhibited substantial differences in their moderation actions, (iv) a remarkable fraction of the database data is inconsistent, (v) the platform X (formerly Twitter) presents the most inconsistencies. Our findings have far-reaching implications for policymakers and scholars across diverse disciplines. They offer guidance for future regulations that cater to the reporting needs of online platforms in general, but also highlight opportunities to improve and refine the database itself.
Introduction
@@
Content moderation refers to the strategies and processes through which an online platform ensures that the content and behaviors it hosts adhere to the platform’s self-defined standards and guidelines, and to legal regulations (Gillespie 2018). Moderation interventions are the specific actions through which platforms enforce such strategies and processes. Currently, a wide array of interventions are adopted, which involve removing content and users (Jhaver et al. 2021), demoting, restricting, or otherwise reducing their visibility (Le Merrer, Morgan, and Trédan 2021; Trujillo and Cresci 2022), showing users warning messages (Katsaros, Yang, and Fratamico 2022) or attaching warning labels to problematic content (Zannettou 2021), and even acting on monetary rewards (Ma and Kou 2022).
@@
However, despite their importance in maintaining safe and healthy online environments, the application of these interventions often sparks concerns. The controversies are multifaceted, such as considering the potential misuse of removals and reduced visibility decisions as a form of censorship that threatens users’ freedom of speech (Myers West 2018). Furthermore, transparency in content moderation is often questioned, as the processes behind moderation decisions —whether executed by black-box AI algorithms or human moderators— are sometimes perceived as opaque, or lacking clarity and consistency (Gillespie 2020). Issues of explainability, fairness, and the susceptibility of automated systems to biases further contribute to the ongoing debate, highlighting the delicate balance that platforms must strike between upholding community standards and legal obligations, and respecting user rights in the digital sphere (Grimmelmann 2015).
@@
The Digital Services Act. In light of the predominant role held by large online platforms in controlling the flow of information and shaping user opinions, the European Union (EU) recently enacted a strategic response to address the intricate challenges of regulating online spaces. This response also meets the growing demand for legislative intervention to reform the existing model centered on platform self-regulation, wherein social media have long been largely autonomous in defining and implementing their own rules and procedures. The Digital Services Act (DSA) is the regulation officially introduced by the EU on October 2022 to modernize its legal framework governing digital services (European Parliament and Council 2022). Its aim is to create a safer and more accountable online environment for users within the EU, by addressing a range of issues related to online platforms, including content moderation, user safety, and the responsibilities of digital service providers (Turillazzi et al. 2023; Church and Pehlivan 2023). The DSA specifically includes several articles with important implications for online content moderation, the most relevant of which are briefly presented in the following.
@@
  • Article 33 designates those platforms with more than 45M users in the EU as either very large online platforms (VLOPs) or very large online search engines (VLOSEs). Due to their strategic role, these must comply with the most stringent rules of the DSA. On April 2023, the EU designated 17 VLOPs and 2 VLOSEs. [^1] https://ec.europa.eu/commission/presscorner/detail/en/IP˙23˙2413
@@
Research Questions and Contributions. Herein, we carry out an in depth quantitative analysis of very large social media moderation practices, as reflected by their SoR submissions to the DSA-TDB. We analyze all 195.61M SoRs submitted by eight platforms during the first 60 days of the DSA-TDB, providing an early yet thorough look into (i) the overall frequency of moderation interventions by very large social media, (ii) their grounds for decision, (iii) the types of applied restrictions, (iv) the types of moderated content, (v) the timeliness of their interventions, and (vi) their use of automation in content moderation. In carrying out these analyses, we specifically focus on answering to the following research questions.
RQ1:
@@
Significance. To the best of our knowledge, this is the first platform-wise comparative analysis of moderation actions in a vast social media ecosystem at a continental scale. As such, the implications of our work are many and various. On the one hand, our results allow drawing insights into the moderation practices of different platforms, for example, by identifying common patterns and striking differences. We also uncover incoherent data submitted by some platforms, which raises concerns on the reliability of some of the information contained in the DSA-TDB. On the other hand, our analysis also sheds light on some of the challenges implied by imposing a unified reporting structure to a large number of diverse platforms. Our findings can thus inform future developments of the DSA-TDB, so as to possibly overcome the limitations of the current structure.
Background
@@
The DSA marks a crucial advancement in EU digital services regulation, aiming to enhance online content moderation and citizen protection. It focuses on improving online safety, transparency, and provider accountability (Wilman 2022; Cauffman and Goanta 2021). By emphasizing user protection from illegal content and ensuring fundamental rights, the DSA enforces stringent measures on strategic platforms within the EU —particularly VLOPs and VLOSEs— to reduce risks and foster a safer online space (Husovec 2024).
Recent literature in the realm of digital services regulation offers insightful perspectives on the DSA and on its implications for online content moderation. For example, Urman and Makhortykh (2023) carried out a detailed analysis of the transparency reports released by large online platforms pursuant to Article 15 of the DSA. They found partial compliance with the Santa Clara Principles 2.0 (Santa Clara Principles 2021) among tech companies, with Snapchat fully compliant and Google excelling in state-involved reporting. Their findings also highlighted widespread non-compliance in flagging processes, underscoring a need for enhanced transparency under the DSA. The works by Nunziato (2023) and Church and Pehlivan (2023) explored the global impact that the DSA had on social media content moderation. By comparing the EU’s approach with the DSA to the US’s emphasis on free speech, they suggest that the stringent requirements of the former could set new global internet content moderation norms. Schneider and Rizoiu (2023) went one step further and assessed the potential for harm reduction caused by the DSA, by means of self-exciting point processes. Their analysis, based on real-world data, indicates the DSA’s capacity to standardize and effectively regulate content moderation, with an emphasis on rapid action against highly harmful content.
@@
Instead of focusing on the broad implications of the DSA as a whole, other scholars investigated its possible effects in more specific contexts. For example, Leerssen (2023) investigated the lawful basis for inherently opaque moderation interventions such as demotions and “shadow banning” (Le Merrer, Morgan, and Trédan 2021), revealing a strong tension between platforms’ moderation needs and the transparency-enhancing framework of the DSA. Finally, Strowel and De Meyere (2023) discussed the implications of the DSA in combating online disinformation. The analysis underscored platform transparency requirements and their potential in curbing misleading content. It also contemplated the challenges and effectiveness of these obligations, proposing enhancements to the regulatory framework of the DSA for improved enforcement and compliance.
Overview of the DSA Transparency Database
@@
The main entity in the DSA-TDB is the statement of reasons (SoR) —specific information regarding the removal or restriction to a certain content by providers of hosting services. SoRs are intended to empower users to understand, and challenge, content moderation decisions. Herein, we analyze all SoRs submitted during the initial 60 days since the establishment of the database, from 25 September to 23 November 2023. The data, documentation, and source code of the DSA-TDB are publicly available under permissive licenses on its official website. [^3] https://transparency.dsa.ec.europa.eu/ For our analyses, we used the “light” version of the daily archives, released as compressed CSV files with 33 fields. With respect to the the full database schema, the “light” version excludes four lengthy fields: three free text attributes and one territorial scope attribute. Based on a brief analysis of these four attributes on a sample of a few days of the “full” archive, we decided for the “light” version, as we considered them uninformative in answering our research questions.
@@
Out of the 17 VLOPs designated by the DSA, 16 have submitted at least one SoR. Wikipedia is the only VLOP that did not make any submission to the DSA-TDB during the initial 60-day period. In this time, a total of 551.06M SoRs were submitted. The largest submitter was by far Google Shopping (48.3%). Given our focus on social media —that is, platforms centered on user generated content and social networking— we selected the following eight platforms of interest for this study: Facebook, Instagram, LinkedIn, Pinterest, Snapchat, TikTok, X (formerly Twitter), and YouTube. With 195.61M SoRs, these platforms represent 35.5% of the whole database.
@@
In addition to the DSA-TDB, we also took into account data available on the most recent DSA Transparency Reports released by the eight selected social media (see Table 3 in the Appendix). These compulsory reports mainly cover the first semester of the year (H1 2023) and were all made publicly available around October on the websites of the respective platforms. In particular, we manually extracted the average monthly authenticated active recipients (MAAR) of each platform for H1 2023. MAAR are registered and logged-in active recipients of the service. In the context of moderation, the number of MAAR is more relevant than that of all active recipients, because an active recipient that is not authenticated is usually not allowed to create or generate content on a platform, which is in turn the main target of moderation decisions. Moreover, in the absence of detailed data on the total content created for this period, the average MAAR is a useful, if faulty, metric to relativize the number of SoRs by platform. Table 1 presents an overview of these metrics for the eight selected social media.
[…]
@@
Table 1: Overview of very large social media platforms, ordered by the number of statements of reasons (SoRs) per average monthly authenticated active recipients (MAAR).
Analyses and Results

### 20 Jan 2024 v2

@@ Abstract.
Since September 2023, the Digital Services Act (DSA) obliges large online platforms to submit detailed data on each moderation action they take within the European Union (EU) to the DSA Transparency Database. From its inception, this centralized database has sparked scholarly interest as an unprecedented and potentially unique trove of data on real-world online moderation. Here, we thoroughly analyze all 353.12M records submitted by the eight largest social media platforms in the EU during the first 100 days of the database. Specifically, we conduct a platform-wise comparative study of their: volume of moderation actions, grounds for decision, types of applied restrictions, types of moderated content, timeliness in undertaking and submitting moderation actions, and use of automation. Furthermore, we systematically cross-check the contents of the database with the platforms’ own transparency reports. Our analyses reveal that (i) the platforms adhered only in part to the philosophy and structure of the database, (ii) the structure of the database is partially inadequate for the platforms’ reporting needs, (iii) the platforms exhibited substantial differences in their moderation actions, (iv) a remarkable fraction of the database data is inconsistent, (v) the platform X (formerly Twitter) presents the most inconsistencies. Our findings have far-reaching implications for policymakers and scholars across diverse disciplines. They offer guidance for future regulations that cater to the reporting needs of online platforms in general, but also highlight opportunities to improve and refine the database itself.
social media, content moderation, digital services act, online regulation
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; Sometime 01–02, 2024; Neverlandisbn: 978-1-4503-XXXX-X/YY/ZZccs: Human-centered computing Empirical studies in collaborative and social computingccs: Information systems Social networking sitesccs: Social and professional topics Governmental regulations
1. Introduction
@@
Content moderation refers to the strategies and processes through which an online platform ensures that the content and behaviors it hosts adhere to the platform’s self-defined standards and guidelines, and to legal regulations (Gillespie, 2018). Moderation interventions are the specific actions through which platforms enforce such strategies and processes. Currently, a wide array of interventions are adopted, which involve removing content and users (Jhaver et al., 2021), demoting, restricting, or otherwise reducing their visibility (Le Merrer et al., 2021; Trujillo and Cresci, 2022), showing users warning messages (Katsaros et al., 2022) or attaching warning labels to problematic content (Zannettou, 2021), and even acting on monetary rewards (Ma and Kou, 2022).
@@
However, despite their importance in maintaining safe and healthy online environments, the application of these interventions often sparks concerns. The controversies are multifaceted, such as considering the potential misuse of removals and reduced visibility decisions as a form of censorship that threatens users’ freedom of speech (Myers West, 2018). Furthermore, transparency in content moderation is often questioned, as the processes behind moderation decisions —whether executed by black-box AI algorithms or human moderators— are sometimes perceived as opaque, or lacking clarity and consistency (Gillespie, 2020). Issues of explainability, fairness, and the susceptibility of automated systems to biases further contribute to the ongoing debate (Nogara et al., 2023), highlighting the delicate balance that platforms must strike between upholding community standards and legal obligations, and respecting user rights in the digital sphere (Grimmelmann, 2015).
1.1. The Digital Services Act
@@
In light of the predominant role held by large online platforms in controlling the flow of information and shaping user opinions, the European Union (EU) recently enacted a strategic response to address the intricate challenges of regulating online spaces. This response also meets the growing demand for legislative intervention to reform the existing model centered on platform self-regulation, wherein social media have long been largely autonomous in defining and implementing their own rules and procedures. The Digital Services Act (DSA) is the regulation officially introduced by the EU on October 2022 to modernize its legal framework governing digital services (European Parliament and Council, 2022). Its aim is to create a safer and more accountable online environment for users within the EU, by addressing a range of issues related to online platforms, including content moderation, user safety, and the responsibilities of digital service providers (Turillazzi et al., 2023; Church and Pehlivan, 2023). The DSA specifically includes several articles with important implications for online content moderation, the most relevant of which are briefly presented in the following.
@@
  • Article 33 designates those platforms with more than 45M users in the EU as either very large online platforms (VLOPs) or very large online search engines (VLOSEs). Due to their strategic role, these must comply with the most stringent rules of the DSA. On April 2023, the EU designated an initial set of 17 VLOPs and 2 VLOSEs, [^1] https://ec.europa.eu/commission/presscorner/detail/en/IP_23_2413 whereas on December 2023 it identified a second set of 3 VLOPs focused on adult content. [^2] https://ec.europa.eu/commission/presscorner/detail/en/ip_23_6763
@@
Figure 1. Timeline of relevant events concerning the DSA Transparency Database.
1.2. Research Questions and Contributions
Herein, we carry out an in depth quantitative analysis of very large social media moderation practices, as reflected by their SoR submissions to the DSA-TDB. We analyze all 353.12M SoRs submitted by eight platforms during the first 100 days of the DSA-TDB, providing an early yet thorough look into (i) the overall frequency of moderation interventions by very large social media, (ii) their grounds for decision, (iii) the types of applied restrictions, (iv) the types of moderated content, (v) the timeliness of their interventions, and (vi) their use of automation in content moderation. Figure 1 depicts this 100-day period of study (orange-colored) in the context of relevant events related to the DSA-TDB. In carrying out these analyses, we specifically focus on answering to the following research questions.
RQ1::
@@ 1.2.1. Significance
To the best of our knowledge, this is the first in-depth, platform-wise comparative analysis of moderation actions over an extended time period, in a vast social media ecosystem, at a continental scale. As such, the implications of our work are many and various. On the one hand, our results allow drawing insights into the moderation practices of different platforms, for example, by identifying common patterns and striking differences. We also uncover incoherent data submitted by some platforms, which raises concerns on the reliability of some of the information contained in the DSA-TDB. On the other hand, our analysis also sheds light on some of the challenges implied by imposing a unified reporting structure to a large number of diverse platforms. Our findings can thus inform future developments of the DSA-TDB, so as to possibly overcome the limitations of the current structure.
2. Related Work
We first present an overview of the implications of the DSA for online regulation, particularly on content moderation. We then consider the critical aspect of platform transparency for content moderation, both in the form of reports by platforms themselves and in the form of access to platforms’ data for independent analysis. Finally, we discuss a couple of precursory analyses of the DSA-TDB in its earliest stage, and how the present work improves upon these.
2.1. The implications and impact of the DSA
@@
The DSA marks a crucial advancement in EU digital services regulation, aiming to enhance user protection in an ever-increasing digital world. It focuses on improving online safety, transparency, and provider accountability (Wilman, 2022; Cauffman and Goanta, 2021). Moreover, the DSA contains stringent measures —particularly for VLOPs and VLOSEs— to reduce risks and foster safer platforms for millions of users that could inspire other jurisdictions to update their current norms (Husovec, 2024). For these reasons, the DSA has sparked scholarly debate on its potential implications and impact, in Europe and beyond.
For instance,  Nunziato (2023), along with Church and Pehlivan (2023), explored the global impact that the DSA could have on social media content moderation. They suggest that the stringent requirements of the DSA could set global Internet moderation norms due to the high cost of implementing different systems for different jurisdictions. Given the attractiveness and power of the European single market, platforms are likely to use worldwide their already DSA-compliant systems. This is also likely to create some friction with other important markets, such as the United States of America, as the more stringent approach by the EU contrasts the approach by the USA, which often prioritizes freedom of speech. In particular, the pivotal Section 230, enacted in 1996 as part of the USA Communications Decency Act, states that online platforms are not to be treated as publishers, and are thus not liable for content created by users within the platform (230, 1996). Such provision has been both heralded as a saviour of free speech and criticized as a shield for scoundrels (Ardia, 2009). The DSA also safeguards service providers in a similar manner in Articles 14 and 15. However, it is much more stringent on the obligations of service providers and in the kinds of contents deemed illegal and the actions required. In this view, Schneider and Rizoiu (2023) assessed the potential for harm reduction caused by the DSA by means of self-exciting point processes. Their analysis, based on real-world data, indicates that the DSA has the potential to standardize and effectively regulate content moderation, with an emphasis on rapid action against highly harmful content.
@@
Instead of focusing on the broad implications of the DSA as a whole, other scholars investigated its possible effects in more specific contexts. For example, Leerssen (2023) studied the lawful basis for inherently opaque moderation interventions such as demotions and “shadow banning” (Le Merrer et al., 2021), revealing a strong tension between platforms’ moderation needs and the transparency-enhancing framework of the DSA. Finally, Strowel and De Meyere (2023) discussed the implications of the DSA in combating online disinformation, underscoring platform transparency requirements and their potential in curbing misleading content. The work also reflected on the challenges and effectiveness of these obligations, proposing enhancements to the regulatory framework of the DSA regarding enforcement and compliance.
2.2. Platform transparency on content moderation
Social media platforms, crucial as they are in shaping public discourse, face scrutiny over their content moderation practices. Balancing freedom of expression with digital civil discourse is challenging (DeNardis and Hackl, 2015; Gillespie, 2018), and criticism arises due to a perceived lack of transparency and biases in moderation (Jhaver et al., 2019; Myers West, 2018). Far-right users claim censorship (Jahanbakhsh et al., 2021; Myers West, 2018), while left-leaning users point to policy enforcement inconsistencies (Shen and Rosé, 2019). In addition, marginalized communities often face disproportionate moderation (Haimson et al., 2021; Thach et al., 2022). The evolution from manual to automated moderation (Zhang et al., 2021; Annamoradnejad and Habibi, 2023) further complicates matters, with mixed user opinions on their fairness and bias (Lyons et al., 2022; Gonçalves et al., 2023; Nogara et al., 2023). Therefore, transparent content moderation practices by online platforms are of paramount importance. In this regard, the DSA provides three important mechanisms: periodic transparency reports, granting researchers access to platform data, and the new centralized transparency database —the subject of the present study.
Transparency reports are an accountability instrument to detail user and authority notifications, content removals, and accuracy of a platform moderation system. These reports —already mandatory for VLOPs and VLOSEs under the DSA— provide critical insights into digital governance, enhancing public understanding and regulatory oversight (Buri and van Hoboken, 2021; Commission, 2023b). For example, Urman and Makhortykh (2023) carried out a detailed analysis of the transparency reports released by large online platforms pursuant to Article 15 of the DSA. They found partial compliance with the Santa Clara Principles 2.0 (Santa Clara Principles, 2021) on transparency and accountability in content moderation, with Snapchat being fully compliant and Google excelling in state-involved reporting. Their findings also highlighted widespread non-compliance in flagging processes, underscoring a need for enhanced transparency within the DSA.
Having access to data posted on social media through specific APIs or historical repositories is also important toward moderation transparency. Arguably, the then free Twitter Researcher API (Chen et al., 2022) and Reddit Pushshift archive (Baumgartner et al., 2020) were among the most open and widely used data for social media analysis. For several years, these enabled researchers to investigate the effects of moderation policies of their respective platforms. For instance, several studies have investigated the impact of major moderation interventions on user behavior in Reddit, at the community (Trujillo and Cresci, 2022) and individual levels (Trujillo and Cresci, 2023), inside the platform as well as across platforms (Horta Ribeiro et al., 2021). Interestingly, in several cases —an particularly in the case of political communities— such moderation interventions had mostly mixed effects (Chandrasekharan et al., 2022; Shen and Rosé, 2022).
Moderation interventions by Twitter were also found to have mixed effects. For instance, deplatforming controversial influencers was followed by a general decrease in conversations and toxicity, but also by an increase in offensive content among certain users (Jhaver et al., 2021). In the case of Trump’s ban after the Capitol attack, there was even increased polarization and ideological differences (Alizadeh et al., 2022). More recent studies have shown that Twitter was also the target of content manipulation in the context of the Ukraine invasion and the French presidential election in 2022 (Pierri et al., 2023). Similar phenomena has been observed in historically less open platforms, such as Facebook and Instagram, across different topics, ranging from political issues (Kalsnes and Ihlebæk, 2021; Goldstein et al., 2023) to groups focused on unhealthy behaviours (Chancellor et al., 2016; Gerrard, 2018).
In early 2023, due to changes in management and in part to the unauthorized use of social media content to train generative AI models, the Twitter API ceased to be free [^4] https://twitter.com/XDevelopers/status/1621026986784337922 (even for research) and Reddit severely restricted access to its API. [^5] https://redd.it/12qwagm/ Indeed, both platforms, as well as others, changed their terms of service with respect to data access and collection. Consequently, many social media analysis projects were halted, significantly hindering the transparency and accountability of social media. In this respect, the DSA includes a provision for granting vetted researchers access to VLOPs and VLOSEs data, provided that certain conditions are met (European Parliament and Council, 2022). The overarching goal is that of understanding how society is shaped by these platforms, as well as supporting their oversight of DSA compliance. Some of the very large platforms have just recently set up these new mechanisms to access EU content, hence the effect of such provision on online moderation research remains to be seen.
2.3. Precursory analyses of the DSA Transparency Database
The DSA-TDB, an unprecedented potential data trove regarding moderation actions, has piqued much scholarly interest since its announcement. Even at this early stage, two non-peer-reviewed works provided preliminary analyses of the initial data submitted by large social media platforms to the database. The study by Dergacheva et al. (2023) only took into consideration submissions made during the very first day of the database. However, in spite of the extremely limited breadth of the analysis, they found much heterogeneity in the reported degrees of automation, applied visibility measures, and types of content moderated by various online platforms on a single day. In a work more similar to our present contribution, Drolsbach and Pröllochs (2023) examined the first 60 days of data from the database, revealing relevant disparities in moderation practices across different platforms. Their findings underscored the predominance of automated moderation, with X (formerly Twitter) being the main exception, and highlighted some inconsistencies in implementing the DSA guidelines. Differently from our study, the work by Drolsbach and Pröllochs (2023) is mostly descriptive and it does not systematically take into consideration the differences among platforms. For instance, a regression analysis was conducted regarding the use of automated means for moderation, which however conflated different levels of automation and data from various platforms. Given that TikTok represented more than 60% of the data, their resulting model is heavily influenced by the self-reported practices of a single platform, with very different dynamics compared to the other seven social media.
On the other hand, our research presents a different approach to the database and includes several aspects not covered in any of the previous analyses. Firstly, we investigate the initial 100 days of the database, providing a much more comprehensive and robust time frame. Secondly, we conduct a thorough platform-wise analysis of moderation actions, taking into consideration the potential distinctiveness of each platform. Thirdly, our study provides a deeper assessment of the data submitted to the database, specifically addressing questions about the adequacy, consistency, and timeliness of submitted SoRs. Lastly, we also carry out a systematic coherence cross-check of the database against each platforms’ public DSA Transparency Reports, which allows to identify discrepancies and inaccuracies in the reported data. Our approach not only uncovers worrisome inconsistencies, but also critically evaluates the effectiveness and limitations of the current database schema, while suggesting possible enhancements.
3. Overview of the DSA Transparency Database
@@
The main entity in the DSA-TDB is the statement of reasons (SoR) —specific information regarding the removal or restriction to a certain content by providers of hosting services, intended to empower users to understand and challenge content moderation decisions. Herein, we analyze all SoRs submitted during the initial 100 days since the establishment of the database, from 25 September 2023 to 02 January 2024. The data, documentation, and source code of the DSA-TDB are publicly available under permissive licenses on its official website. [^6] https://transparency.dsa.ec.europa.eu/ For our analyses, we used the “light” version of the daily archives, released as compressed CSV files with 33 fields. With respect to the the full database schema, the “light” version excludes four lengthy fields: three free text attributes and one territorial scope attribute. Based on a brief analysis of these four attributes on a sample of a few days of the “full” archive, we decided for the “light” version, as we considered them uninformative in answering our research questions.
@@
Out of the 17 VLOPs initially designated by the DSA, 16 have submitted at least one SoR. Wikipedia is the only VLOP that did not make any submission to the DSA-TDB during the initial 100-day period. In this time, a total of 1.4861B SoRs were submitted. The largest submitter was by far Google Shopping, with more than a billion SoRs (67.8% of the total), and the smallest was Zalando, with merely 9 SoRs submitted. Given our focus on social media —platforms centered on user generated content and social networking— we selected the following eight platforms of interest for this study: Facebook, Instagram, LinkedIn, Pinterest, Snapchat, TikTok, X (formerly Twitter), and YouTube. With 353.12M SoRs, these platforms account for 23.8% of the whole database.
@@
In addition to the DSA-TDB, we also took into account data available on the most recent DSA Transparency Reports released by the eight selected social media (see Table 3 in appendix A). These compulsory reports mainly cover the first semester of the year (H1 2023) and were all made publicly available around October 2023 on the websites of the respective platforms. In particular, we manually extracted the average monthly authenticated active recipients (MAAR) of each platform for H1 2023. MAAR are registered and logged-in active recipients of the service. In the context of moderation, the number of MAAR is more relevant than that of all active recipients, because an active recipient that is not authenticated is usually not allowed to create or generate content on a platform, which is in turn the main target of moderation decisions. Moreover, in the absence of detailed data on the total content created for this period, the average MAAR is a useful, if faulty, metric to relativize the number of SoRs by platform. Table 1 presents an overview of these metrics for the eight selected social media.
@@
Table 1. Overview of very large social media platforms, ordered by the number of statements of reasons (SoRs) per average monthly authenticated active recipients (MAAR).
Platform #SoRs %SoRs #MAAR #$\frac{\text{\#SoRs}}{\text{\#MAAR}}$
TikTok $184.772$M $52.33\%$ $135.9$M $1.360$
Pinterest $60.297$M $17.08\%$ $124.0$M $0.486$
Facebook $79.050$M $22.39\%$ $259.0$M $0.305$
YouTube $19.267$M $5.46\%$ $445.8$M $0.043$
Instagram $8.111$M $2.30\%$ $259.0$M $0.031$
Snapchat $1.119$M $0.32\%$ $102.0$M $0.011$
X $0.466$M $0.13\%$ $61.3$M $0.008$
LinkedIn $0.038$M $0.01\%$ $45.2$M $0.001$
4. Analyses and Results